# X-Dancer: Expressive Music to Human Dance Video Generation

### v1

@@ 4.1 Implementation Details
Dataset. Our model is trained on a curated visual-audio dataset comprising 107,546 monocular recordings of human dance performances in both indoor and outdoor settings, with an average clip duration of 30 seconds. Each video is cropped to a resolution of $896\times 512$ and resampled to $30$ fps, encompassing a broad spectrum of half- to full-body dances across diverse music genres and showcasing individuals with varied body shapes and appearances.
Training. We train our full pipeline in three stages. First, we train a multi-part pose VQ-VAE to encode and quantize 60 joint coordinates, with keypoint confidence scores, into 5-part pose tokens. The pose of each body part uses 6 tokens, each with a unique 512-entry codebook of 6-dimensional embeddings. This VQ-VAE is trained for 40k steps with a batch size of 2048 at a learning rate of $2\times 10^{-4}$. Next, we train an autoregressive model for pose token prediction with 300k steps, initialized with pretrained GPT-2 model weights, using a batch size of 24 and a learning rate of $1\times 10^{-4}$. The model is trained with 64-frame pose sequences with a context window of 2224 tokens in total, enabling extended dance motion generation via sliding segments during inference. In the final diffusion stage, we fine-tune the denoising UNet and ReferenceNet with two randomly selected frames from a video, followed by co-training the motion guidance decoder and the temporal module with diffusion losses on consecutive 16 frames. We train the diffusion stage with 90k steps at a learning rate of $1\times 10^{-5}$ and a batch size of 16. All stages are trained with 8 A100 GPUs using the AdamW optimizer [44].
Inference. For inference, we initiate our autoregressive dance motion generation from pose tokens encoded from the reference image pose, maintaining skeleton consistency with the specified body shape. Extended dance sequences are generated in 64-frame sliding segments with a 12-frame overlap, while 8 uniformly sampled frames from previous generated motions serve as global motion context. Using the full generated pose token sequences, we synthesize all video frames simultaneously, applying prompt traveling [36] to improve temporal smoothness.
@@
4.2 Evaluations and Comparisons
To the best of our knowledge, no existing approach addresses music-driven 2D video generation from a single human image. For baseline comparisons, we adapted and combined established models to create two baselines. First, we adapted the audio-driven, diffusion-based portrait animation model Hallo [42], retraining it for our task by substituting its audio features with our music embeddings to animate human images via cross-attention modules. For the second baseline, we utilize 3D music-to-dance generation models like Bailando [32] and EDGE [37] for motion synthesis, projecting their outputs into 2D skeleton maps, which are then fed into a diffusion model with a pose guider for controlled image animation. We note that during our submission, the AIST dataset [38] was unavailable due to maintenance, and we were unable to train our model on the same dataset as Bailando [32] and EDGE [37]. We conduct separate evaluations of all models on the test set of AIST++ [21] (40 videos) as well as on a test split of our curated music-video dataset (230 videos).
@@ 4.2 Evaluations and Comparisons
Quantitative Evaluation. We numerically compare our method with baselines in terms of quality of both motion generation and video synthesis. Specifically, we calculate the Fréchet Video Distance (FVD) [39, 10] between generated dance motions and ground-truth training sequences for assessment of motion fidelity. For motion diversity, we compute the distributional spread of generated pose features (DIV) [21, 32], for which the generated motion should aim to match the scores of the ground truth distribution rather than maximize their absolute values. To numerically evaluate the alignment between the music and generated dancing poses, we follow [21, 32] to measure the Beat Align Score (BAS) by the average temporal distance between each music beat and its closest dancing beat. These evaluations are conducted in 2D pose space, where we project the 3D skeleton motion of Bailando and EDGE into 2D and detect poses over synthesized videos from the retrained Hallo model. As shown in Tab. 2, our method surpasses all baselines in terms of motion fidelity (FVD), while achieving the second-best music beat alignment (BAS) and diversity (DIV). Notably, EDGE is trained on sequences from professional dancers, whereas our dataset comprises videos of everyday individuals, reflected in the beat alignment score of the ground-truth videos. Despite this, our method significantly outperforms Bailando trained on AIST++ and Hallo trained on our curated dataset for beat alignment. Hallo entangles motion generation and video synthesis, and achieves a higher DIV score than X-Dancer primarily due to its extremely noisy video outputs which results in jittering and chaotic skeleton motions following pose projection.
@@
For evaluation of video synthesis fidelity, we measure the FVD and FID-VID [3] score between the ground-truth and generated dance videos. Additionally, we assess identity preservation using the ArcFace score [8], which measures the cosine similarity of face identity features (ID-SIM). All metrics are evaluated on our test video dataset. As an extra baseline, we replace the motion generator in EDGE and Bailando with our trained transformer model and generate the animation using a pose guider. As shown in Tab. 2, our method achieves the highest visual quality and identity preservation, which we attribute to our disentangled design of motion generation and video synthesis (compared to Hallo) and the use of an implicit motion token decoder rather than an explicit pose guider.
@@
Qualitative Comparisons. We present visual comparisons between our method and the baselines in Fig. 3. For more dynamic comparisons and user study on perceptual quality, please refer to our supplementary material. The modified Hallo [42] represents an end-to-end diffusion pipeline that directly synthesizes the final video without intermediate motion generation. However, it exhibits noticeable artifacts, particularly in large articulated body motions, and often fails to preserve the human body’s articulation topology. Bilando [32] and EDGE [37] generate body motion in 3D space without considering the scene context or the human shape in the reference image. Despite post-processing for skeleton alignment in the pose guider input, these methods still struggle with significant identity and shape distortions, often producing unnatural interactions with the background scene. Furthermore, they do not model head and hand motions, leading to more rigid and less expressive dance movements compared to X-Dancer.
Finetuning for Characterized Choreography. While our method operates as a zero-shot pipeline, generalizing seamlessly to new reference images and music inputs, it can also be fine-tuned for characterized choreography using only a few sample dance videos. This adaptability is challenging for 3D motion generation models like EDGE [37] and Bailando [32], which require extensive effort in creating 3D dance movements. As shown in Fig. 4 and our supplementary video, our method successfully captures and mimics the specific choreography after fine-tuning with only 30 dance videos from diverse performers, showcasing its efficiency and versatility in adapting to specific dance styles.
@@ 4.3 Ablation Study
We conduct ablation studies by systematically removing individual design components from our full training pipeline. First, we assess the effectiveness of the multi-part VQ-VAE in capturing subtle and expressive human movements. Compared to a single-part whole-body VQ-VAE, our approach achieves a significantly lower L1 reconstruction error per joint (0.5px vs. 0.83 px), with pronounced improvements in the hands (0.4px vs. 0.64px) and head (0.42px vs. 0.52px). This improvement translates into superior motion diversity, expressiveness, and control precision in the diffusion model, as shown through quantitative metrics and visual examples in our supplementary material. Next, we assess the impact of global music and motion context on motion generation. As presented in Tab. 3, both contexts contribute to producing more consistent, plausible, and music-synchronized motions. We further analyze the benefits of 2D motion modeling and transformer-based autoregressive generation by scaling both model parameters and dataset size. Across all metrics, we observe significant performance gains (Tab. 3) as the number of training parameters increases from 117M (GPT-small) to 345M (GPT-medium) and data scale from 10k to 100k videos, underscoring the scalability potential of monocular dance video data with our architecture, sheding light on further performance improvements as they scale. In Tab. 2, we compare our token-based motion guidance with AdaIN against a pose guider with explicitly decoded skeleton map. Our method demonstrates reduced motion jittering and enhanced temporal consistency due to the temporal motion context, and showcases superior identity and body shape preservation, as also visually shown in our supplementary video.
@@ 5 Conclusion
Limitations and Future Work. Our model is trained solely on curated real-human daily dance videos from internet, which can be noisy and the dance movements may not exhibit the precision found in professional dancer videos. Consequently, out-of-domain human images may lead to rendering artifacts, and the generated dance motions may occasionally lack alignment with music. While we designed our pipeline to be end-to-end trainable and scalable, we currently implement it in stages due to memory limitations. Future work will explore large-scale, multi-machine training to further enhance performance and efficiency.

### v2

@@ 4.1 Implementation Details
Dataset. Our model is trained on a curated in-house visual-audio dataset of 76,818 monocular dance recordings in indoor and outdoor settings, averaging 15 seconds per clip. Each video is cropped to $896\times 512$ resolution and resampled at 30 fps, covering half- to full-body dances across diverse music and performer characteristics. Details on the dataset are provided in the supplementary paper.
Training. We train our full pipeline in three stages on 8 A100 GPUs using the AdamW optimizer [57]. First, we train a multi-part pose VQ-VAE to encode and quantize 60 joint coordinates and keypoint confidences, into 5-part pose tokens. Each body part’s pose uses 6 tokens, each with a unique 512-entry codebook of 6D embeddings. This VQ-VAE is trained for 40k steps with a batch size of 2048 at a learning rate of $2\times 10^{-4}$. Next, we train an autoregressive model for pose token prediction over 300k steps, initialized with pretrained GPT-2 weights, using a batch size of 24 and a learning rate of $1\times 10^{-4}$. The model operates on 64-frame pose sequences with a 2224-token context window. Lastly, we fine-tune the denoising UNet from SD1.5 [37] and ReferenceNet with two randomly selected frames per video, followed by co-training the motion token decoder and temporal module [14] with diffusion losses on consecutive 16 frames. The diffusion stage is trained for 90k steps at a learning rate of $1\times 10^{-5}$ and a batch size of 16.
Inference. We initiate our autoregressive dance motion generation from pose tokens encoded from the reference image pose. Extended dance sequences are then generated in 64-frame sliding segments with a 12-frame overlap, while 8 uniformly sampled frames from last segment serve as global motion context. We synthesize all video frames simultaneously with the full pose token sequence, applying prompt traveling [43] to improve temporal smoothness.
4.2 Evaluations and Comparisons
To the best of our knowledge, no existing approach addresses music-driven 2D video generation from a single human image. We adapted and combined established models to create two baselines for comparisons. First, we adapted the audio-driven, diffusion-based portrait animation model Hallo [53], retraining it for our task by substituting its audio features with our music embeddings to animate human images via cross-attention modules. For the second baseline, we utilize 3D music-to-dance generation models like Bailando [38] and EDGE [44] for motion synthesis, projecting their outputs into 2D pose maps, which are then fed into a diffusion model with a pose guider for controlled image animation. For fair comparison, we also train our motion transformer (stage 2, X-Dancer-AIST) on the AIST dataset [45], consistent with Bailando and EDGE, but using detected 2D poses instead of 3D skeletons. However, since AIST has limited diversity in identities and appearances, we use our diffusion model—trained on our curated dataset—for comparison on synthesized videos. We conduct separate evaluations of all models on the test split of AIST (40 videos) as well as our curated music-video dataset (230 videos).
@@ 4.2 Evaluations and Comparisons
Quantitative Evaluation. We numerically compare X-Dancer with baselines in terms of quality of both motion generation and video synthesis. Specifically, we calculate the Fréchet Video Distance (FVD) [46, 12] between generated dance motions and ground-truth training sequences for assessment of motion fidelity. For motion diversity, we compute the distributional spread of generated pose features (DIV) [25, 38], To numerically evaluate the alignment between the music and generated dancing poses, we follow [25, 38] to measure the Beat Align Score (BAS) by the average temporal distance between each music beat and its closest dancing beat. These evaluations are conducted in 2D pose space, where we detect poses over synthesized videos from the retrained Hallo model, and project the 3D skeleton motion of Bailando and EDGE into 2D via orthographic projection, using camera parameters that best align with the human bounding box in the reference image.
As shown in Tab. 2, our method surpasses all baselines in terms of motion fidelity (FVD) and music beat alignment (BAS), while achieving the second-best diversity (DIV). Notably, Bailando and EDGE is trained on professional dances (AIST++ [25]), whereas our dataset comprises videos of everyday individuals, reflected in lower BAS even for ground-truth videos. Our method trained on AIST (X-Dancer-AIST) achieves the highest BAS whereas our model trained on our curated videos still significantly outperforms Bailando and Hallo for beat alignment. Hallo entangles motion generation and video synthesis, and achieves a higher DIV score than X-Dancer primarily due to its extremely noisy video outputs which results in jittering and chaotic skeleton motions following pose projection.
Table 4: Ablation on VQ-VAE and video diffusion. PG and MD denotes PoseGuider and Motion Decoder respectively.
Pose L1 FullBody/Head/Hands $\downarrow$
Single-Part VQVAE 0.83 / 0.64 / 0.52
Multi-Part VQVAE 0.50 / 0.40 / 0.42
Video PSNR $\uparrow$ / LPSIS $\downarrow$ / FVD $\downarrow$
GT Pose + PG 19.465 / 0.197 / 294.52
Single-Part + MD 17.079 / 0.258 / 350.73
Multi-Part + PG 18.836 / 0.207 / 384.16
Multi-Part + MD 19.148 / 0.207 / 295.87
@@
For evaluation of video synthesis fidelity, we measure the FVD and FID-VID [4] score between the ground-truth and generated dance videos. Additionally, we assess identity preservation using the ArcFace score [10], which measures the cosine similarity of face identity features (ID-SIM). All metrics are evaluated on our test video dataset. As an extra baseline, we replace the motion token decoder in our full pipeline with a pose guider (Our motion + PG). As shown in Tab. 2, our method achieves the highest visual quality and identity preservation, which we attribute to our disentangled design of motion generation and video synthesis (compared to Hallo) and the use of an implicit motion token decoder rather than an explicit pose guider.
@@
Qualitative Comparisons. We present visual comparisons between our method and the baselines in Fig. 3. For more dynamic comparisons and user study, please refer to our supplementary material. The modified Hallo [53] represents an end-to-end diffusion pipeline that directly synthesizes the final video without intermediate motion generation. However, it exhibits noticeable artifacts, particularly in large articulated body motions, and often fails to preserve the human body’s articulation topology. Bilando [38] and EDGE [44] generate body motion in 3D space without considering the scene context or the human shape in the reference image. Despite post-processing for skeleton alignment in the pose guider input, these methods still struggle with significant identity and shape distortions, often producing unnatural interactions with the background scene. Furthermore, they do not model head and hand motions, leading to more rigid and less expressive dance movements compared to X-Dancer. We also note that our model trained on monocular dance videos is efficient and versatile, easily adapting to specific dances. This is not easily achievable with EDGE or Bailando, which require extensive effort in crafting 3D skeleton dances (supplementary paper Sec.3)
@@ 4.3 Ablation Study
We conduct ablation studies by systematically removing individual design components from our full training pipeline.
Multi-part VQ-VAE. As shown in Tab. 4, our compositional multi-part VQ-VAE significantly reduces pose reconstruction error compared to single-part whole-body pose tokenization. The multi-part VQ-VAE better captures fine-scale pose variations, while single-part tokenization often loses high-frequency details, resulting in more rigid and less expressive motion generation. To further assess its impact, we train a video generation diffusion model using single-part pose tokenization. As illustrated in Figure 4, the single-part representation struggles to control fine-grained motions like hip swaying, leg lifting and head tilting, leading to perceptually lower-quality videos. This is further supported by its inferior video reconstruction metrics (PSNR and LPIPS [64]) in Tab. 4.
Figure 4: Our Multi-part VQ-VAE captures intricate poses such as hand gesture (left), arm bending, head tilting and leg lifting(right) whereas single-part pose tokenization fails to fully replicate them.
Motion Generation. Next, we assess the impact of global music and motion context on motion generation. As presented in Tab. 3, both contexts contribute to producing more consistent, plausible, and music-synchronized motions. We further analyze the benefits of 2D motion modeling and transformer-based autoregressive generation by scaling both model parameters and dataset size. Across all metrics, we observe significant performance gains (Tab. 3) as the number of training parameters increases from 117M (GPT-small) to 345M (GPT-medium) and data scale from 10k to 100k videos, underscoring the scalability potential of monocular dance video data with our architecture, sheding light on further performance improvements as they scale.
Pose-Guided Video Synthesis. In Tab. 2, we compare our motion token decoder (X-Dancer) against a pose guider using explicitly decoded skeleton map (Our motion + PG). While our transformer-generated motion exhibits jittering, our motion token decoder significantly reduces jitter and enhances temporal consistency by leveraging temporal motion context. Additionally, it demonstrates superior identity and body shape preservation compared to the pose guider.
In Tab.4, we evaluate the effectiveness of our motion token decoder (Multi-Part + MD) against the pose guider (Multi-Part + PG) in self-driven video synthesis. Our motion decoder directly infers body pose coordinates and confidences from pose tokens, achieving lower reconstruction errors (PSNR and LPIPS) compared to the pose guider, which relies on explicitly decoded pose maps. However, due to motion blurriness in fast dance motions and fine-grained structural variations in small pixel regions like the face and hands, visual rendering artifacts remain present in both methods. Notably, these artifacts stem primarily from the diffusion model’s limitations rather than errors in pose reconstruction, as evidenced by the similar video error observed in GT Pose+PG (Tab. 4).
@@ 5 Conclusion
Limitations and Future Work. Our model is trained solely on curated real-human daily dance videos, which can be noisy and lack the motion precision found in professional dancer videos. Consequently, out-of-domain human images may lead to rendering artifacts, and the generated dance motions may occasionally lack music alignment. More failure cases are present in the supplementary paper. While we designed our pipeline to be end-to-end trainable and scalable, we currently implement it in stages due to memory limitations. Future work will explore large-scale, multi-machine training to further enhance performance and efficiency.
@@
Supplementary Material
In the supplementary material, we provide additional details on our curated in-house dataset (Section 6). In Section 7, we provide a user study to further compare X-Dancer to all baselines. We discuss more use cases of  X-Dancer in Section 8, and present some failure cases and discuss some limitations of our method (Section 9). For more dynamic visual results, please refer to our offline webpage.
6 Dataset
Our in-house video dataset is sourced from a third-party curation service, featuring monocular recordings of everyday dance performances by diverse individuals worldwide. To ensure quality, we filter out videos shorter than 5 seconds, those with low resolution or poor quality (HyperIQA score [40] below 40), videos containing multiple people or with a human bounding box smaller than $27.5\%$ of the frame, and footage with non-static cameras (corner pixel standard deviation exceeding 20). After filtering, the dataset comprises 76,818 dance videos, totaling around 360 hours. All filtered videos are center-cropped to $896\times 512$ resolution and resampled to 30 fps.
To analyze the dataset, we employ the state-of-the-art vision-language model (VLM) Qwen2.5-VL [3], yielding the following statistics:
  • Gender distribution: $78\%$ female, $22\%$ male.
  • Age distribution: $68\%$ of dancers are classified as young, with the rest as middle-aged or older.
  • Ethnicity distribution: 49% White, 26% Asian, 11.2% Latino, 9% Black, with the rest others.
  • Video duration distribution: 10.04% of videos are between 5-10 seconds, 33.06% between 10-15 seconds, 49.72% between 15-20 seconds, 6.25% between 20-25 seconds, and 0.94% exceed 25 seconds.
  • Recording environment: 88% of videos are recorded indoors.
  • Dance styles: 89% are categorized as freestyle/hip-hop, while the remaining consist of popping and locking.
  • Motion characteristics: 52% feature medium-level movements (e.g., hip swaying, arm waving), while 47.4% exhibit strong motion variations (e.g., leg lifts, body rotations, and translations). 96% of videos do not contain large body turning.
Different from the AIST [45] and FineDance [26] dataset which capture professional dancers in a multiview setup, our dataset consists of monocular recordings of everyday individuals, offering wider accessibility and greater diversity in both dance motions and identity features. However, compared to professional dancers, our curated performances may exhibit weaker beat alignment, mostly frontal movements and less distinct genre characteristics, often leaning towards freestyle movement.
7 User Study
In addition to our quantitative and qualitative comparisons to various baseline methods provided in Section 4.2 of our main paper, including Hallo [53], Bailando [38] + PoseGudier, and EDGE [44] + PoseGuider. We conduct a user study to further assess perceptual quality, motion fidelity and consistency with the music across all methods.
We generated 25 dance videos for each method conditioned on randomly selected reference images with music tracks. Each video is 10 seconds long, recorded at 30 FPS, totaling 300 frames per video. Evaluation was conducted using a questionnaire distributed via Google Sheets. Participants were asked to choose the best-performing video among candidates based on five specific metrics: (1) Human Identity Consistency, (2) Music Beat Alignment, (3) Music Style Alignment, (4) Motion Consistency and Naturalness, and (5) Overall Quality. In total, 375 responses were gathered from 15 independent participants with no prior knowledge of video generation techniques. For each question sample, the videos from different methods were randomly permuted. The reference images were also provided to assist participants in assessing identity preservation.
Our method substantially surpasses all the baselines across different metrics, as evidenced in Table 5.
Table 5: User study. The values indicate user preference ratios (%).
Method ID Consistency Beat Alignment Style Alignment Motion Consistency Overall
Hallo [53] 18.67 2.13 1.60 1.87 1.87
Bailando [38]+PG 3.73 10.93 6.40 3.73 3.47
EDGE [44] + PG 4.80 24.80 17.87 11.47 9.87
X-Dancer 72.80 52.13 74.13 82.93 84.80
8 More Visual Results.
Please refer to our webpage for all the dynamic visual results.
Single Reference, Multiple Music. Given a single reference image, X-Dancer demonstrates the ability to generate diverse and expressive dance motions, maintaining consistent movement styles that adapt to different music genres and beat flows. This underscores X-Dancer’s capability to effectively interpret the global music context while synchronizing seamlessly with local rhythmic beats.
Single Music, Multiple References. We present diverse dance videos generated by X-Dancer for various reference images, all driven by the same music track. While maintaining a consistent dance style synchronized to the shared music, each generated video also reflects the personalized attributes derived from its corresponding reference image, showcasing X-Dancer’s adaptability and attention to individual identity details.
Single Music, Single Reference. We demonstrate a variety of dance movements generated from a single reference image, all driven by the same music track. While all dance movements are well-aligned with the music beats, our model exhibits the ability to produce diverse and dynamic dance motions, highlighting its versatility and creativity.
Finetuning for Characterized Choreography. While our method operates as a zero-shot pipeline, generalizing seamlessly to new reference images and music inputs, it can also be fine-tuned for characterized choreography using only a few sample dance videos. This adaptability is challenging for 3D motion generation models like EDGE [44] and Bailando [38], which require intricate multiview captures or extensive effort in creating 3D dance movements. As shown in Fig. 5 and our supplementary video, our method successfully captures and mimics the specific choreography after fine-tuning with only 30 dance videos from diverse performers, showcasing its efficiency and versatility in adapting to specific dance styles.
Additional Results. We provide additional results including baseline comparisons and results of X-Dancer-AIST (discussed in line 449 - 455 of the main paper).
9 Limitations and Failure Cases.
In addition to the limitations and future work discussed in Section 5 of our main paper, we would like to discuss some additional limitations and failure cases. Specifically, noticeable rendering artifacts remain observable, particularly in the face and hands, and color flickering or over-saturation can occasionally occur. However, as indicated in Table 4 of the main paper (GT Pose + PG), these artifacts persist even when using ground-truth poses, and the overall video quality improvement remains marginal. Such issues are challenges across human image animation models and are orthogonal to the core contributions of our work . Future work could explore stronger video diffusion base models such as [34, 56, 23], and incorporate specialized embeddings or priors for these fine-grained regions [33] to mitigate these problems. Furthermore, while our method effectively generates motions informed by the identity and shape of the reference image, there are cases where the generated motions become inconsistent with the background scene or reference identity. This misalignment can lead to unnatural motions and, in severe cases, pronounced rendering artifacts (Figure 6).
Figure 6: Failure cases. Please note the rendering artifacts on the face and hands, as well as the unnatural renderings under challenging poses.